a)     A hypothetical economy is given by the following identities:

C = 3000

I = 2000

G = 2500

T = 0.2Y

MPC = 0.5

X=6500

Z=5500 + 0.2Y

                   i.           Find the equilibrium level of income.

               


1
Expert's answer
2021-07-05T16:24:17-0400

"Y=3,000+0.5(Y\u22120.2Y)+2,000+2,500+6,500\u22125,500\u22120.2Y"


"Y=3,000+0.4Y+2,000+2,500+6,500\u22125,500\u22120.2Y"

"Y=14,000\u22125,500+0.2Y"


"Y=8,500+0.2Y"

"0.8Y=8,500"

"Y=10,625"

Equilibrium level of income "= 10625"